OpenStudy (anonymous):

FACTOR COMPLETLY 49X^2+25=

OpenStudy (ash2326):

(7x)^2+5^2=0 (7x)^2=-25 (7x)=-5i or +5i so x=-5i/7 and +5i/7 i= root of -1